A man borrow Tk.60,000 from another man at the rate of 6 percent compound interest on condition that




he will repay the money in 8 equal installments. What will be the size of the installment?

1
Expert's answer
2022-01-03T16:13:00-0500

V=Tk.60,0000, r"=6\%" ,n=8

"V=\frac{A}{r}[1-(1+r)^{-n}]"

"Tk.60,000=\frac{A}{0.06}[1-(1+0.06)^{-8}]"

"Tk.60,000=\frac{A}{0.06}[1-0.63]"

"Tk.60,000×0.06=0.37A\\A=\frac{3600}{0.37}=9729.8\approx Tk.9730"